Reports 1-1 of 1 Clear search Modify search
ryutaro.takahashi - 8:35 Saturday 17 April 2021 (16499) Print this report
Comment to Height adjustment of ITMX (16421)

(Log on 16th)

[Takahashi, Sato, Ikeda]

The adjustment of height with pistons was canceled. The BS FR has reached the upper limit (picture). The keystone height of BF was set to +/-0mm.

Images attached to this comment
Search Help